OpenStudy (anonymous):

Georgia is purchasing treats for her classmates. Georgia can spend exactly $10 to purchase 25 fruit bars, each equal in price. Georgia can also spend exactly $10 to purchase 40 granola bars, each equal in price. The equation is .40x+.25y=10.00. I just need to know how to graph it.

OpenStudy (bibby):

you can get it into y = mx+b form by subtracting .40 x

OpenStudy (anonymous):

So you would come out with .25y=-.40x+10

OpenStudy (bibby):

Then you have .25y=-0.40x+10.00 Multiply each term by 4 to get: y=-1.6x+40.00

OpenStudy (anonymous):

So on the graph you would go up 40 then go down 1.6 over 1

OpenStudy (bibby):

1.6 over 1? Even if you didn't multiply by 4, you'd make a table and plot some points

OpenStudy (anonymous):

What? I'm confused...

OpenStudy (anonymous):

Why wouldn't go down 1.6 over 1?

OpenStudy (bibby):

You have 2 lines .25y=-0.40x+10.00 y=-1.6x+40.00 they're the same exact line

OpenStudy (anonymous):

I have to graph it.

OpenStudy (bibby):

so make a table.

OpenStudy (anonymous):

So if you buy 2 fruit bars you'll be able to buy 36.8 granola bars?

OpenStudy (bibby):

Where'd you get those numbers from? 2 fruit bars cost 0.80. That means there's 9.20 left for granola bars.

OpenStudy (anonymous):

I'm literally so confused. Will you make the table?

OpenStudy (bibby):

oh, you were right

OpenStudy (bibby):

lol, sorry. I'm a tad braindead right now.

OpenStudy (anonymous):

Im still confused.

OpenStudy (bibby):

y represents the fruits bars, x represents the granola bar

OpenStudy (anonymous):

I know that

OpenStudy (the_fizicx99):

What is you need bibbeh :o how to graph it?

OpenStudy (bibby):

yeah graph it pls

OpenStudy (anonymous):

Yesss

OpenStudy (the_fizicx99):

Gimmie the equation .-.

OpenStudy (anonymous):

.40x+.25y=10

OpenStudy (bibby):

it's up in the first post. .40x+.25y=10.00

OpenStudy (the_fizicx99):

.40x+.25y=10 -.40x -.40x 0.25y = -0.40x + 10 /0.25 /0.25 /0.25 y = -1.6x + 40 plot it at (0,40) and apply rise / over it's negative so go down 1.6 and run 1

OpenStudy (the_fizicx99):

http://prntscr.com/3aoi8r

OpenStudy (bibby):

forgive me :<

OpenStudy (the_fizicx99):

I forgive you :*

OpenStudy (anonymous):

So what's the slope of the line?

OpenStudy (the_fizicx99):

-1.6

OpenStudy (bibby):

y = -1.6x + 40 is in the form y = mx + b where m is the slope

OpenStudy (the_fizicx99):

Go bibbeh ~ ♦o♦

OpenStudy (anonymous):

Okay and explain what the slope means in the context of Georgia purchasing treats?

OpenStudy (the_fizicx99):

Idk about that, I only came here to fix the equation :> bib you got this ;)

OpenStudy (anonymous):

ughh

OpenStudy (bibby):

it represents the amount of granola bars